### v4

Abstract
Intrusion Detection Systems (IDS) are widely employed to detect and mitigate external network security events. Vehicle ad-hoc Networks (VANETs) are evolving, especially with the development of Connected Autonomous Vehicles (CAVs). It is, therefore, crucial to assess how traditional IDS approaches can be utilised for these emerging technologies. We combat this issue by introducing a stacked ensemble learning scheme for IDS to improve the detection of threats. Our method employs several machine learning (ML) algorithms to identify threats more precisely than single methods. We demonstrate our approach’s performance compared to current ML methods on CICIDS2017 and VeReMi and establish that it can better discriminate threats. Our process also incorporates hyperparameter optimisation and feature selection to improve performance further. Overall, our results suggest that stacked ensemble learning is a promising technique for enhancing the effectiveness of IDS.
keywords:
[…]
1 Introduction
Vehicles are increasingly being connected to their operating environments, resulting in Vehicle Ad-hoc Networks being considered a subset of Mobile Ad-hoc Networks when conditions are altered and environmental changes occur due to the continual rotation of automobiles. Inter-vehicle communication is a growing research, standardisation, and development area as more cars are configured with wireless communication and computing technology. The new era of automobiles and road traffic systems will depend on wireless networks and the associated data to better coordinate and manage traffic flow safely and effectively through developments such as machine learning models. Ultimately, the innovations of VANETs must enable safer and more efficient travel for road users. To enable this innovation, we need a mechanism by which vehicles can be made sufficiently intelligent to manage their process for road safety. This notion constituted the foundation for VANETs. Software-defined networking (SDN) based VANETs follow the below architecture Ku et al. (2014). VANET is a wireless network in which both stationary and mobile nodes connect. Such a system has frequent disconnections, a dynamic topology, and significant node mobility. Each attribute may result in a security flaw that hazards the system. To detect known and unencrypted attacks in a local detection scenario, researchers are showing great interest in current years due to the SDN concept in VANET that implements the ML-based detection mechanism. IDS is still in its infancy, and many studies must be conducted to make ID as successful as possible. Current IDS is plagued by difficulties and obstacles that require rapid and intensive investigation. This work has identified significant elements and activities that research communities must address. The problem and assignment areas :
  • Incomplete data set: A data set is a collection of comprehensive data and information from a survey that must be examined. From IDSs, data sets contribute significantly to the accuracy of outcomes. For example, researchers are now utilising DARPA 98, 99, and the University of New Mexico’s immune system data sets. Still, we cannot prevent relatively recent attacks due to their lifetime.
  • Methodologies for detection: It is a crucial aspect of determining if a packet/data originates from an attack or whether it is valuable data that the user needs for carrying out operations or activities. Detection algorithms must be efficient enough to match all cases in a short amount of time and efficiently match criteria. The basis for the search policy might be abuse or abnormality. Therefore, anomaly or abuse might serve as the detection approach. A declaration is a piece of data utilised by the matching process. A downside is that there are no instructions for matching fresh assaults. Therefore, this attack is detected by determining if a new attack has been discovered or by detecting a change in data that matches the pattern. Thus, we need a good and quick algorithm that will identify patterns wider and faster to fit the majority of assaults.
  • Weak design: All IDSs have a compact design; thus, if a user needs to edit a portion of the IDS, the IDS must be stopped, modified, and redeployed.
By referring to the literature review below, several essential study gaps in ML-based IDS research have been noted. Still, there is enough opportunity to improve the accuracy of the intrusion detection model. Moreover, IDS users still struggle to detect known unencrypted attacks because IDS has not yet achieved 100% accuracy. One significant issue with IDS is that they regularly alert the auditor to false positives. However, an IDS can be optimised to reduce the number of false positives and false negatives, which is a deficiency of publicly available VANET databases with attacks. No fast, inexpensive, and effective in-built privacy-preserving ML-based IDS is still visible.
1.1 Contributions
This article proposes a new model for anomaly detection in the SDN-based VANET for the Binary and Multiclass classification to detect different attacks and Benign in the VeReMi dataset. The contributions of this paper are as follows:
  1. 1.
    This research suggests a novel approach to intrusion detection in SDN-based VANETs that integrates different machine-learning approaches for better accuracy and speed of operation.
  2. 2.
    Furthermore, the suggested IDS model can identify minor attack entries in a big dataset, an essential component of VANET security.
  3. 3.
    The evaluation of our suggested approach using two comprehensive datasets (CICIDS2017 and VeReMi) can provide a more comprehensive evaluation of IDS models in SDN-based VANETs. VeReMi offers a large and diverse network traffic data set, while CICIDS2017 focuses on intrusion detection in a simulated environment relevant to VANET security. This combination can help evaluate IDS performance in detecting and responding to a broader range of security threats. Insights into the strengths and weaknesses of the proposed ML model and techniques can also be gained.
2 Background
VANET, a type of wireless network connecting both stationary and mobile nodes, is characterised by frequent disconnections, a dynamic topology, and high mobility, which can lead to potential security flaws. Compared to Mobile Ad Hoc Networks (MANET), VANETs exhibit considerably greater mobility and more stringent topology restrictions. Vehicular Ad Hoc Networks (VANETs) enable applications such as real-time traffic monitoring, blind crossing prevention, dynamic route planning, collision prevention, and internet access for vehicles. According to the World Health Organization (WHO), 1.3 million people die in road crashes annually, with 20 to 50 million suffering non-fatal injuries WHO (2023).
Figure 1: Software-defined VANET
The new era of automobiles and road traffic systems will depend on wireless networks and machine learning. As stated previously, many lives have been lost on highways. Consequently, when road traffic increases, road safety must also grow. We need a mechanism by which vehicles can be made sufficiently intelligent to manage their process for road safety. This notion constituted the foundation for VANETs. SDN-based VANETs follow the below architecture Ku et al. (2014).
2.1 Communication And Standerd
The USDOT (The United States Department of Transportation) has designated Dedicated Short-Range Communication as the preferred inter-vehicle communication medium. The communication infrastructure for linked automobiles is provided via WAVE protocols. IEEE 802.11p protocol and IEEE 1609.x protocols define WAVE. Their activities conform to the internet protocol stack. Table I displays WAVE’s tiered architecture. Physical and data connection layers are supported by IEEE 802.11p. The tiers of operation for IEEE 1609.x are network, transport, and application. This protocol focuses on the physical and Media Access Control (MAC) levels to provide low latency and good dependability across short radio communication lines. Chekkouri et al. (2015) It functions at both the physical and MAC levels. It offers a direct wireless interface and mapping between MAC and physical data. In addition, it provides media access rules and aids in multi-channel operations at the MAC layer.
2.1.1 IEEE 802.11p
enables security via authentication and encryption of private communications. All cars are expected to be registered by a Motor Vehicle registration unit. This would assign MAC IDs to automobiles. Alternatively, Vehicle Identification Numbers (VINs) can serve as MAC addresses for automobiles. Vehicles will obtain IP addresses upon network connection. Vehicle identification will be based on their MAC addresses, IP addresses, or publicly provided IDs. (Zheng and Wu, 2016)
2.1.2 IEEE 1609.4p
Support for multi-channel operations is achieved by time division so radios can switch between different channel intervals. Control Channel (CCH) and Service Channel (SCH) are the two channels of this IEEE standard. Continuous safety messages are sent via CCH, whilst other applications are routed via SCH. IEEE 1609.3p defines the network and transport layer services. Here, IP addressing and routing occur. IEEE 1609.2p is the protocol that specifies secure message exchanges and encryption. IEEE 1609.1p is a protocol specifying alert and probe vehicle message forms. Currently, a modification to IEEE 1609.x is in the works. The following enhancements have been made. Draft IEEE 1609.5 to handle vehicle-to-vehicle and vehicle-to-infrastructure connectivity. Draft IEEE 1609.6 to provide remote management between Onboard Units and Road-Side-Units. IEEE Standard 1609.11 for safe electronic transactions.IEEE Standard 1609.12 defines the allocation of value identifiers in WAVE. (Chen, Jiang and Delgrossi, 2009)
2.2 SDN strategies to address VANET security concerns
SDN can manage network-wide transmissions without configuring devices separately, providing a holistic view and centralised control. These features enhance the security of Vehicular Ad Hoc Networks (VANETs) by defending against common vulnerabilities. Akhunzada et al. (2016). Nevertheless, the network will be most susceptible to security concerns if a rogue node compromises or impersonates the controller. An attacker may issue commands to network devices to undertake malicious actions through the hacked controller. Through programmability, attackers can build false and inaccurate flow rules to interrupt communication and manipulate devices and software. These harmful operations are capable of causing network resource unavailability, accidental occurrences, traffic congestion, and communication delays. The following section discusses various risks associated with SDN-based VANET. These two conceptions are at odds with whether SDN is a security benefit for VANET or a security risk for VANET.
2.3 VANET Characteristics and Negatives
VANET is a wireless network in which both stationary and mobile nodes connect. Such a system has frequent disconnections, a dynamic topology, and significant node mobility. Each attribute may result in a security flaw that hazards the system. Due to the high mobility of nodes, VANET topology is dynamic and unpredictable; this and other circumstances, such as severe weather conditions, may cause VANET to have several disconnections throughout the system.
2.4 Intrusion Detection System (IDS)
The intrusion detection system is a network-based system that detects any harmful activity on the target network. The software architecture operation depends on observing the network and noting any attempts or incidences of network infiltrations into the system as the necessary actions are done in time to prevent the systems from getting harmed. When an IDS identifies malicious behaviour in a network or node, it also blocks the malicious node’s access to the network to avoid additional network damage. Intrusion Detection System designs can be centralised or decentralised. A centralised intrusion detection system collects and analyses data in a centralised system, whereas a distributed intrusion detection system collects and analyses data on several dispersed hosts. A centralised IDS achieves effective intrusion detection for a single system. The sites for data analysis remain constant regardless of the number of observed hosts. In situations involving numerous systems, IDS employs a decentralised design. In decentralised intrusion detection systems, numerous data security analyses are performed at several network locations. The number of places where data analysis is conducted is proportional to the number of observed host computers. This is the practical system architecture achieved by evaluating the network and detecting any intrusion attempts or malicious activities that made the network vulnerable so that appropriate measures could be taken promptly to prevent such activities from causing damage. There are five distinct categories of VANET attacks: (1) authenticity, (2) confidentiality, (3) availability, (4) integrity, (5) accountability, and (6) privacy. The Intrusion Detection System counteracts harmful actions by blocking a questionable IP address, port, or person from accessing and gaining access to the network. Depending on the goal of intrusion detection systems, they are separated into three types.
A network-based intrusion detection system (NIDS) operates at the network level to identify any rogue network node or incursion attempt in the network. NIDS are strategically positioned within the network to manage traffic to and from all equipment. It analyses passing traffic on the whole subnet and compares it to a library of known assaults. The administrator can be notified once an attack or suspicious behaviour is detected. Host-based Intrusion Detection Systems (HIDS) function locally on a particular node to detect intrusions. HIDS operate on particular network hosts or devices. A HIDS analyses only the inbound and outbound packets from the device and alerts the user or admin if any unusual behaviour is noticed. It compares the current snapshot of system files to the past snapshot. If vital system files are updated or deleted, the administrator is notified to examine them. Wireless Intrusion Detection Systems (WIDS) operate wireless networks with functionality comparable to NIDS. Based on their detecting strategies, IDS is categorised under Misuse Detection and Anomaly Detection and, based on the behaviour of IDS.
3 Related Works
As a part of the research, different articles have been reviewed to explain how earlier researchers have ended work on an Intrusion Detection System for cyber-attack detection in SDN-based VANET. This Literature review study employs the methods stated by Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA). The methodology covers eight main steps: User needs assessment, define the Research Questions, Search String, Selection of Source, Select Relevant Paper, Screening Findings, Extract Data, and finally, synthesise the collected data. (Kitchenham and Charters, 2007)
3.1 User Needs Assessment
To effectively determine user requirements for cyber-attack protection, the following essential steps should be taken:
3.1.1 Step 1: Identify Stakeholders
Government and private organisations are the two main stakeholders in the cybersecurity field. In the private sector, companies from various sectors of the economy, the ISPs, software vendors, end users etc.
Table 1: CICIDS2017 Dataset attacks.
Attacks Size
DoS 251712
PortScan 158804
DDoS 128025
Brute Force 13832
Web Attack 2180
Bot 1956
Infiltration 36
Heartbleed 11
Table 2: VeReMi Dataset attacks and size.
Attacks Size
BENIGN 437429
Attack type 16 (Eventual stop Attack) 56595
Attack type 4 (Random Attack) 30510
Attack type 2 (Constant Offset Attack) 30473
Attack type 1 (Constant Attack) 30473
Attack type 8 (Random Offset Attack) 29460
3.1.2 Step 2: Understand IDS Issues
IDS and Malware detection systems can still not detect unknown, Encrypted, and Inference attacks with 100% accuracy. According to Kaspersky, most Malware detection products do not even reach the required 90% security threshold. A significant player in the IDS market, CISCO, claimed that Rapid detection—those made in less than five minutes—and long-term learning capacities displayed excellent performance in their IDS. Cisco Encrypted Traffic Analytics, or IDS, discovered nearly two-thirds of all known malicious flows in less than five minutes. Moreover, Cisco declared this IDS demonstrated better identification results even with low volumes of 0 to 20 flows, whereas 2000 plus flows enabled 100% detection for known attacks. Nevertheless, this solution has a limitation with a small attack size and taking three hours to gain 100% accuracy. However, the Cisco Adaptive Security Device Manager (ASDM) to manage Cisco Secure Firewall ASA (with IDS/IPS) has been impacted by Common Vulnerabilities and Exposures CVE-2022-20829, CVE-2021-1585, and CVE-2022-20828. Researcher Jake Baines, who disclosed them to Cisco in February and March of 2022, added that many users were undoubtedly not updating their Cisco firewalls properly. CISCO (2022) Scroxton (2022).
3.2 Intrusion detection system challenges in Software-defined vehicular network (SDVN)
There are several difficulties in creating an adaptable and versatile IDS using ML in SDN-based VANET networks due to the need for effective data classification in academic research. To have a consistent and accurate evaluation of intrusion detection systems, it is essential to construct robust datasets. For testing and evaluation of intrusion detection, the following datasets are already available:
3.2.1 CICIDS2017
The CICIDS2017 dataset offers a benchmark for evaluating intrusion detection systems with various network characteristics and attack scenarios, including Brute Force, Heartbleed, Botnet, DoS, DDoS, Web Attack, and infiltration. The dataset collected over five days in July includes over 80 traffic characteristics per flow, obtained using the CICFlowMeter tool from the Canadian Institute for Cybersecurity. However, this dataset may not cover all authentic network behaviours due to individual differences in attackers’ strategies and timings. Sharafaldin et al. (2018). Nevertheless, CICIDS2017 holds more than 280,000 entries with all unfilled features. Very few numbers of entries contain ’NaN’ or ’Infinity.’
3.2.2 VeReMi
The VeReMi dataset is a synthetic benchmark dataset for misbehaviour detection systems in Vehicular Ad Hoc Networks that relies on the VEINS framework addendum. VeReMi is implemented with the VEINS framework, modified from Version 4.6, with the help of LuST (Luxembourg Simulation of Urban Mobility Traffic). This dataset can be efficiently employed to test the performance of misbehaviour detection approaches implemented on a city-wide scale while utilising fewer computing resources than standard (frequent) VEINS simulations. The dataset comprises On-Board Unit (OBU) message logs, including timestamps, sender IDs, positions, speeds, and reception times. VeReMi also includes a verification file, which provides GUI parameters on the validity of the data of each message upon which the verification confirms whether it is a "normal" or "attack" case. van der Heijden et al. (2018)
The dataset spreads across three attacker densities, five distinctive attack types, and three levels of density; it also includes various scenarios at different times of the day, including how traffic varies across the time of the day with the vehicle density range from 35 vehicles in low-dense areas to 519 vehicles during maximum congestion; and diverse types of attacks where the attacks in this paper are attacks against the position expressed in many ways from the position remaining still to noising out the position. While versatile, the VeReMi dataset is restricted in terms of both position-falsifying attacks and all types of network threats and advanced malicious strategies. Being a synthesised dataset, it lacks real-life cases but can be developed further by researchers.
The dataset is publicly available for reproducibility; thus, the research results made from the dataset are reliable and verifiable. VeReMi has 562,467 entries, out of which 437,429 entries are benign and 125,038 are attacking entries. The dataset comprehensively covers a city-level context of various types of roads and traffic behaviours, thus making it a reliable framework for benchmarking and testing the performance of misbehaviour detection approaches in VANETs. van der Heijden (2018)
3.2.3 Attack detection techniques
The article discusses the ML-based tactic to detect DDoS Attacks in SDVN by Singh et al. (2018). This paper suggests a method for spotting threats in SDN-based vehicular networks based on machine learning. The research has a particular emphasis on vehicle-to-infrastructure (V2I) links. It analyses the current environment to establish the ML strategy that will most successfully detect threats posed by distributed denial of service (DDoS) attacks (Short-lived Transmission Control Protocol Flooding, User Datagram Protocol Flooding). To accomplish this, it collects seven different elements of actual data from TCP (Transmission Control Protocol) and UDP (User Datagram Protocol) flows both when DDoS attacks them and when they are not. The network is educated by employing several different supervised learning algorithms depending on the data acquired, and it is discovered that the gradient boost classifier produces the most satisfactory outcomes. In addition, the random forest, decision tree, and linear support vector machine (linear-SVM) approach all produce results that are close to being the best. On the other hand, the limited training data could be to blame for the weak performance of Neural Networks (NN). Accurate, measurable data was not found in the write-up or the chart.
Aneja et al. (2018) presents an AI-based system to identify a Route Request (RREQ) flooding attack. Researchers used a genetic algorithm (GA) to optimise feature subset selection and an Artificial Neural Network (ANN) for classification. The model scored 99% accuracy from Network Simulator 2 (NS2) 1000 entries simulated dataset.
Yu et al. (2018) In this study, they analyse DDoS assaults in an SDN-based vehicular network by employing TCP flood, UDP flood, or ICMP (internet control message protocol) flood. In this research paper, the authors present a scheme for detecting Distributed Denial-of-Service (DDoS) attacks in SDVN atmospheres using three separate models. A "trigger detection model" for inbound packets and another method is the "flow table feature-based detection model." Both methods use the OpenFlow protocol features and the detection of a threat model based on support vector machine classification. All of these models are used in conjunction with one another. They synthesised and simulated network traffic to accommodate various DDoS attacks via Scapy and hping3, and the findings suggested an efficiency of over 97%. In the article, ML-Based Intrusion Detection Method for VANETs.
Zeng et al. (2018) the authors developed an ML-based Senior2Local IDS with propagation to identify potential threats in VANETs both Globally and Locally. They used a fuzzification ANN technique to recognise malicious attacks globally and protect Cluster Heads in Roadside Units (RSUs). After that, a game theory trust system or RSU is developed separately. They employed a simple and straightforward Support Vector Machine (SVM) to detect attacks locally. The findings demonstrated that the method was more reliable and resilient than earlier ML-based tactics.
Gad et al. (2021) conduct the experiment on a ToN-IoT dataset and test various ML approaches in binary and multiclass classification problems in VANET to detect intrusion. On the other hand, LR is used for solving binary classification problems, while multiclass issues are addressed with a one-vs-rest approach.
Liu et al. (2014) suggests employing data mining methods in VANETs to identify existing assaults and uncover previously unidentified attacks. The proposed approach makes three primary contributions. First, it develops a decentralised vehicle network to enable scalable communication. Second, it uses two data mining models to demonstrate the viability of an IDS in VANETs. Lastly, it discovers novel patterns of previously unidentified threats.
The suggested solution uses a cell grid to organise the previously mentioned network. Every cell has a transmission tower, enabling users to communicate with others and the internet. Each will use data mining models and rules to identify new assaults. Therefore, the IDS can develop new rules that may be communicated for each subnetwork. The trucks and the tower cell collect the data sent throughout the network. The authors use Naive Bayes and Logistic Regression classifiers to analyse the obtained data. Finally, the authors use simulation to evaluate their strategy’s performance over a network.
First, the Simulation of Urban Mobility (SUMO) simulator produces a file that tracks movement. Then, to mimic a wireless network, this information is given into ns-2. This scenario includes 150 cars that turn randomly at junctions, drive at the posted speed limit (which ranges from 5 to 20 meters per second), and stop at traffic lights that are randomly positioned. Five automobiles were outfitted with Linux to validate the IDS and made to run various network applications. After that, nine months were spent using TCPdump to generate the dataset. There were four different kinds of attacks, with 39 different types documented. After that, WEKA is used to categorise them. Finally, the metrics recall, F-measure, and Matthews’s correlation coefficient were utilised to assess the models.
3.2.4 Each method has yet to reach optimum accuracy
Türkoğlu et al. (2022), With the help of feature selection and hyperparameter optimisation, the DT classifier attained an accuracy score of 99.35%, which was the best possible result. The excellent results achieved with the methodology developed proved the considerable significance of feature selection and optimisation techniques. These findings were demonstrated by the outstanding results obtained with the method.
Kim et al. (2017) proves the existence of a collaborative security threat detection method within software-defined vehicular cloud architecture. A traditional vehicular environment has some exceptional features, such as high flexibility, dynamic networking, and rapid connectivity. His proposed model said a Speed Dependent Volume Controller (SDVC) had collected the flow information from the vehicular cloud (VC) to train multiclass SVM. In the next step, the SVM classifier identifies the bad actors among the network’s nodes. However, performance is still under 90% for Accuracy score, Precision, and Recall.
However, Misra et al. (2011) offers an LA-based system for IDS in VANET. The suggested methodology protects users’ anonymity by giving each car a unique, changing ID. The proposed approach involves developing a VANET administration system. There is a base station at the beginning and end of each route, and all cars have transmitters to stay in touch with the base stations. To redirect traffic, each attacker will generate malicious packets. The budgets of the attackers and the system are similar. The more the attacker’s resources, the greater the number of packets it can create. All that is needed to identify the attackers is their dynamically assigned identifiers. If unique IDs are assigned to vehicles, and many vehicles have the same ID, then at least one of the vehicles is likely malevolent. Under attack, grid budgets may be attributed, and sampling rates can be recalculated with the help of learning automata. Simulations run on their simulator are used to evaluate models. The success rate for detecting four attackers ranges from around 44% to 77%. In the suggested setup, the IDS is located in each base station and has complete sight of overall network traffic.
Tian et al. (2010) proposes a Bus Ad-Hoc Network (BUSNet) based IDS tailored to protect VANETs from malicious activity. This hierarchical anomaly detection technique is built on the virtual setup of the mobile bus agents. In this configuration, the buses serve as the cluster heads, collecting and relaying data packets broadcast by passing vehicles through accessing points installed along the roadways. After that, an NN-based classification system takes that data and uses it to spot DDoS assaults. The proposed method is tested in an artificial network constructed with the ns-2 network simulator. The researchers do not disclose the method or traffic simulator utilised to generate the simulation results. Fifty unique vehicles are simulated, each communicating using a constant bit rate (CBR) and 512-byte packet size. During each attack, two vehicles work together to send out updates at small durations of 0.01 seconds at four different checkpoints over 10 seconds. Both vehicles are used in every attack. The authors establish their benchmark by picking a value below which they treat everything as an assault. This number might range from 0.05 to 0.7, which helps to explain the discrepancy in these results. A threshold value of 0.2 is recommended as the best option. Overall, accuracy has not achieved the milestone to protect VANET.
Mejri and Ben-Othman (2014), The proposed fix is a detection system that uses a statistical method, linear regression, and a watchdog to discover greedy behaviour in the MAC layer passively. A watchdog monitors the degree to which active nodes’ access periods correlate as part of the suggested approach. The network is under assault if the correlation coefficient is close to 1 but not precisely 1, and the linear regression straight-line slope is also not close to 1. This can happen when the correlation coefficient is near one but not quite one or when the slope is very close to 1 but not quite 1. The developed program does this by monitoring three metrics: the interval between transmissions, the total time spent transferring data, and the number of connect attempts made by a node. Then, with the aid of the trace file created by SUMO, the solutions’ functioning is analysed in real-time. -3. The SUMO scenario is based on a city map with labels and signals. Researchers first simulate the network using nodes that behave correctly to ensure the linear regression method is reliable. After that, four greedy automobiles are produced by injecting greedy nodes one at a time. To detect greedy behaviour, this approach takes 1.3 seconds for 1 node, 1.9 seconds for 2 nodes, 3.1 seconds for 3 nodes, and 7.9 seconds for 4 nodes. Ali Alheeti et al. (2015) The solution proposed in the paper is an ANN-based ill-use and anomaly intrusion detection system that can identify DoS assaults. The authors designed the solution by first developing a mobility scenario within SUMO. For ns-2 to read the files SUMO created, a conversion was performed using Mobility Vehicle (MOVE). In the end, ns-2 was utilised to imitate the vehicles’ communication ability. When considering the mobility and traffic situation, the authors used the urban mobility of Manhattan as their inspiration. The simulation takes two minutes and fifty seconds, and it involves cars and six Road roadside units (RSUs). A CBR (Constant Bit Rate) traffic that conveys User Datagram Protocol (UDP) packets is being performed by the vehicles. Only one of the automobiles intentionally tries to cause harm. It will discard the packets rather than send them to the destination. The solution built was accurate 85.02% of the time when classifying normal behaviours and 98.45% when classifying deviant behaviours.
Gruebler et al. (2015) The authors advocate using a Proportional Overlapping Scores (POS) approach to reduce the number of features taken from the trace file. These are fed into an artificial neural network (ANN) to teach it to classify data. After that, the ns-2 trace files that were created were employed in the POS algorithms so that the appropriate characteristics could be extracted. Next, Fuzzification is applied to the dataset’s data to prevent classification difficulties. Finally, the dataset is classified with the help of a Feed-Forward Neural Network (FFNN) using the IDS created for this solution. Sixty thousand records from the dataset were used, with half designated for training, another quarter for testing, and the remaining quarter for validation (25%). This dataset includes both benign and harmful activity in equal measure. The malicious behaviour vehicle was designed as a Black Hole attack, in which the malicious vehicle would not forward any received packets but dump them instead. Ultimately, the intrusion detection system was evaluated using anomaly and misuse detection. Regarding identifying misuse, it acquired a classification of 99.89 % for normal conduct and 99.80% for deviant behaviour. Regarding detecting anomalies, the findings for typical and aberrant behaviours came in at 99.87 and 99.72%, respectively.
Sedjelmaci and Senouci (2015) This paper presents a cluster-based intrusion detection system (IDS) to defend a network against selective forwarding, black holes, wormholes, packet repetition, resource depletion, and Sybil attacks. The technique that has been presented uses several different detection managers that operate on three different levels: cluster associate, cluster head, and RSU. The cars’ speeds determine how they should be organised into clusters. Cluster connection and the assurance of data safety are the criteria utilised in selecting the cluster heads. When social behaviour is introduced, the connectivity inside a cluster is improved. The IDS architecture comprises two primary detection schemes and a decision scheme. Local Intrusion Detection System and Global Intrusion Detection System are the detection systems run at cluster members and heads. The name of the decision-making scheme is Global Decision System, which is hosted on the RSUs. Because of this, the system can monitor several organisations and identify threats on multiple levels. In addition, each level can perform a unique set of algorithms and detecting approaches, which evaluate a unique set of characteristics. The Global Decision System will provide the collective standing of each car, which the cluster head will then send. The Worldwide Decision Scheme will then calculate the trust level of every vehicle. To evaluate the proposed system, ns-3 was used as the network simulator, while SUMO was used to mimic the vehicle’s movement. The IDS was evaluated concerning the detection rate, the false positive rate, and the detection time. The various assaults were put through their paces using a range of vehicle counts during testing. The percentage of hostile cars always equals 45% of the total. The provided findings range from 92% to 100%, although this depends on the total number of vehicles and the type of attack conducted.
Ali Alheeti et al. (2016) proposed A FFNN and SVM-based intrusion detection system is suggested in this paper. A methodical reaction is also meant to protect automobiles if harmful conduct is identified. A dataset composed of SUMO and ns-2 is utilised to train the IDS, and the features extracted from the trace file are used as training data. The POS algorithm has reduced these characteristics from 21 to 15. The collection consists of 30,000 entries that differentiate between appropriate and inappropriate behaviours. Fuzzification is performed on this dataset before it is used for classification. The writers separate the dataset into three categories: validation, test, and training. Generating grey-hole attacks involves picking malicious vehicles that drop packets at arbitrary intervals throughout the attack. The Ad Hoc On-Demand Distance Vector (AODV) protocol needs to be modified so that rushing assault can be carried out successfully. In SVM, the accuracy values achieved through simulation were 99.93% for normal behaviour and 99.64% for aberrant conduct. When FFNN was used, the findings showed an accuracy of 99.82% for normal behaviour and 98.86% accuracy for deviant conduct, respectively.
Table 3: Review of different ML-based IDS models in VANET and SDVN.
Reference ML Models Dataset Attack Featuring Method Top Result
Singh et al. (2018)
GB, RF, DT,
Linear SVM,
LR, NN,
Neural Net.
Generated
Short-lived
TCP Flood,
UDP Flood.
Not measurable
from the chart
Aneja et al. (2018)
ANN
NS2 Trace
RREQ Flood GA based
Accuracy: 95%,
Precision: 97%,
F-Score: 98%,
FPR: 3%
Yu et al. (2018)
SVM
DARPA 2000,
CAIDA
DDoS 2007
TCP flood,
UDP flood,
ICMP Flood,
Etc.
Flow-based
features
and statistical
based entropy
TCP Traffic
Pre-Selection:
DR 98.26 %,
FR 0.52%
After-Selection:
DR 98.56%
FR 0.32%
Zeng et al. (2019)
DeepVCM,
KNN, DT,
1D-CNN, LSTM
ISCX2012,
NS3 Trace
Brute Force,
SSH DDoS,
HttpDoS
RAW data
NS3 Dataset/
DeepVCM/DoS
Precision 98.5%
Recall 98.3%
F1-Score 98.4%
Zeng et al. (2018)
ANN Generate -
CNN
algorithm
applied to
Extract features.
Accuracy:
SMV-CASE 98.7%,
CEAP 98.9%,
Senior2Local 98.37%
Gad et al. (2021)
LR, NB,
DT, SVM,
kNN, RF,
AdaBoost,
XGBoost
ToN-IoT
DoS
DDoS
etc.
Features
Selection:
a. Apply Chi²
b. Apply
SMOTE
c. Apply above
both
With All Features
XGBoost:
Accuracy 99.1%,
Precision 98.4%,
Recall 99.1%,
F1-Score 98.7%
Türkoğlu et al. (2022)
kNN, SVM
DT
Generated DDoS
MRMR
filter
method
DT Algorithm with
25 Features:
Accuracy 99.35%,
Sensitivity 99.22%,
Specificity 99.80%,
F1-Score 99.21%
Kim et al. (2017)
SVM KDD CUP 1999
DoS, Probing,
U2R, R2L
Etc.
Under Diff. Attack
(Approx. from
Graph when $\alpha$=0):
Accuracy
SVM-VC 68.00%,
SVM Nearest
Neighbour 57.00%,
SVM Individual 39.50%
Liu et al. (2014)
NB, LR TCPdump
DoS, R2L,
U2R, Probing
Not measurable
from Chart.
LR showed the
best result.
Ali Alheeti et al. (2015)
ANN NS2 Trace file DoS - Accuracy 85.02%
Gruebler et al. (2015)
ANN
NS2 Trace file and
Animator (32000)
Black Hole - Accuracy 85.02%
Ali Alheeti et al. (2016)
FFNN,
SVM
NS2 Trace file
Grey Hole,
Rushing
Accuracy:
SVM-Abnormal
99.80%
Sharma and Kaul (2018)
SVM NS2 Trace file
Wormhole,
Selective Forwarding,
Packet Drop
Hybrid
Fuzzy
Multi-
Criteria
Feature
Selection
Performance
evaluated
by DR, FPR, DT
Kumar et al. (2015)
Learning
automata
Generated Flooding, Blackhole - -
Shu et al. (2021)
MLP (ANN),
BiGAN
KDD99,
NSL-KDD
DoS, U2R,
R2L, Probing etc.
Accuracy 98.4%
Precision 95.23%
Recall 96.74%
F1-score 95.98%
Precision and recall
are low due to
complex model.
Bangui et al. (2022)
Hybrid model
based on RF
KDD99,
CICIDS2017
DoS, DDoS,
HeartBleed Etc.
- Accuracy 96.93%
Polat et al. (2020)
Softmax
Generate
DDoS - Accuracy 96.9%
Alsarhan et al. (2021)
SVM, GA,
PSO, ACO
NSL-KDD
DoS, Probing,
Unauthorized
access
- Accuracy 98%
Gao et al. (2019)
RF
NSL-KDD,
UNSW-NB15
DDoS - Accuracy 99.95%
Alladi et al. (2021)
CNN-LSTM,
CNN- bidirectional
LSTM, LSTM,
Auto-encoders
Stacked LSTM,
GRU
VeReMi
Various types
of Anomalies
- Accuracy 98%
Alheeti et al. (2017) Using Linear and Quadratic Discriminant Analysis, the authors present an intelligent IDS to guard against assaults, namely DoS and Black Hole attacks. To begin, evil actions are modelled and then emulated. Next, the authors make several adjustments to the AODV protocol to produce the DoS attack. In this scenario, denial-of-service attacks are carried out by dropping packets. After that, multiple mobilities and traffic setups were developed. SUMO and MOVE produced a realistic setting of both abnormal and typical behaviour. In addition to that, ns-2 was utilised to imitate various forms of communication. The dataset is derived from the ns-2 trace file, and all 21 characteristics are preserved throughout the process. The data is given a fuzzy appearance before being put through the test and the training. The findings of the Linear Discriminant Analysis indicate a detection rate of 86.44%, whereas the Quadratic Discriminant Analysis results indicate a detection rate of 85.67%.
Sharma and Kaul (2018) The authors suggest a novel technique for selecting a steady Cluster Head, and it uses Hybrid Fuzzy Multi-standards. After that, an ML-based intrusion detection system that uses SVM is utilised to identify malicious activities. The utilisation of a Dolphin Swarm Algorithm results in an improvement to the SVM-based IDS detection capabilities. This program uses dolphin swarms’ hunting and preying behaviour to locate and isolate misbehaving nodes in a network. The future method is evaluated through the use of simulation. Matlab and NetSim simulate the network, while SUMO models the traffic. During the simulation, a range of node densities from 50 to 300 are examined, with a maximum of 45% of the cars acting as attackers. The suggested intrusion detection system has a detection rate greater than 98%, selective forwarding, and wormhole attack; however, this percentage can vary based on the number of vehicles.
Kumar et al. (2015) offered a cloud intrusion detection system founded on clustering, and Learning Automate (LA) was presented. This method can potentially improve detection activity. The LA determines the leadership of clusters based on factors such as the total comparative velocity and the connectedness grade. The next stage, safeguarding the data, will begin after this procedure. In this process stage, an automaton will validate messages using the HMAC algorithm. When utilising the method that has been suggested, around 93% of harmful actions are recognised, and when mobility is added to leadership formation, the percentage of false positives is reduced. In addition to that, this technique is flexible enough to adapt to any alterations made to the network’s nodes.
Shu et al. (2021) presented an IDS for VANETs by implementing a Software-Defined Networking (SDN)-Based controller on each base station to differentiate between regular network traffic and attack network flows. Using the total network flow data, they trained several SDN controllers for the full VANET using GAN. This IDS approach permits distributed SDN controllers to detect their sub-network flows directly, and it can reduce high processing overheads. The author tested their algorithm using criteria along with precision, Accuracy, F1-Score, recall, and area under the curve (AUC), where maximum accuracy is 98.4%.
Zeng et al. (2019) Researchers have proposed a deep learning-based intrusion detection method incorporating CNN and LSTM. The proposed model’s efficiency was compared to previously developed ML methods using a public dataset and an NS3 simulated VANET dataset with 28236 entries. The recommended model earned an F1 score ranging from 0.91 to 0.99.
Before analysing the above literature, it is essential to consider the findings reported by the National Institute of Standards and Technology (NIST) in 2007: history proved that high rates of false positives and negatives are allied with network-based IDPSs. NIDPSs cannot identify threats inside encoded network traffic; consequently, the system must be installed where the system can screen traffic before encoding or after decoding, or host-based IDPSs must be employed on endpoints to monitor unencrypted movement. Furthermore, NIDPSs are frequently incapable of performing complete analysis under high loads. Furthermore, it requires enough tuning to get good detection accuracy. Scarfone and Mell (2007). In 2022, NIDSs improved to analyse under full load when passing high traffic flow due to modern hardware specs and ML algorithms being used effectively. Besides, researchers should conduct their experiments with an extensive database to get at least 100% accuracy for known attacks. The literature review notes two types of IDS research: A) The ML-based model is built with performance analysis. And B) Innovated Model implementation in a simulator to make a VANET environment to check feasibility before real-time practice.
Gad et al. (2021) discussed ML-based IDS techniques in VANET that are impressive. Researchers used SMOTE, chi2, and SMOTE+chi2 to eliminate the database imbalance problem in the feature engineering phase, and the XGBoost method’s overall score is better than others. In the Training-Testing phase, researchers used Logistic regression (LR), naive Bayes (NB), decision tree (DT), support vector machine (SVM), k-nearest neighbour (kNN), Random Forest (RF), AdaBoost, and XGBoost. This experiment could be better by using the Ensemble Sacking classification to increase overall performance, which this paper wants to address.
Türkoğlu et al. (2022) used the MRMR feature selection algorithm to get the most correlated features for attack detection. First, scholars split the dataset in 3 ways: Training, Validation, and Test Data. After that, researchers used Bayesian Hyperparameter Optimization with training and validation data. However, 3-way splitting (60%, 20% and 20%) was not mandatory for tuning. Instead, the same training data could be used for tuning purposes. That way, 70% or 80% of data could be used for training the database, and 30% or 20% of data will be used for testing purposes to increase the overall accuracy.
Aneja et al. (2018) done proper research using ANN (feed-forward network) with backpropagation technique. This backpropagation operates by scheming the gradient at every layer, known as the local gradient. Then, the weights and bias values are updated according to Levenberg-Marquardt optimisation. Moreover, a Genetic algorithm is used for feature selection.
Alsarhan et al. (2021) use an optimised Support Vector Machine (SVM)-based Intrusion Detection System (IDS) for Vehicular Ad Hoc Networks (VANETs), employing Genetic Algorithm (GA), Particle Swarm Optimisation (PSO) and Ant Colony Optimization (ACO) to enhance accuracy. GA outperformed other algorithms, demonstrating improved detection rates and minimised false positives. A comprehensive evaluation validates the approach on the NSL-KDD dataset, providing significant insights into VANET security.
Figure 2: Binary Classification.
Figure 3: Multiclass Classification (1-Vs-1).
4 Technical Implementation
The VeReMi dataset was created to test the effectiveness of VANET misbehaviour detection systems (vehicular networks). This database contains onboard unit message logs created from a simulation environment and tagged ground truth. The collection contains malicious messages that are meant to cause erroneous application behaviour, which misbehaviour detection techniques are designed to stop. The initial dataset includes five types of position falsification attacks. van der Heijden (2018). This data was generated by van der Heijden (2018), and It has been collected from Sharma (2021).
The CICIDS2017 intrusion detection dataset was used in this study because it comprises the most recent network attack scenarios, including those that are often employed in VANETs. Furthermore, no real security dataset with recent VANET threats is publicly accessible. In contrast, consider the fact that most passenger vehicles are linked to networks for entertainment as well as for transportation. To ensure a good user experience, IDS must identify malware activity from regular network traffic. The benign and common assaults in the CICIDS2017 dataset closely mirror actual real-world data (PCAPs). It also holds the CICFlowMeter network traffic investigation outcomes, which categorises flows based on the time brand, source and endpoint IP addresses, source and endpoint ports, protocols, and attack. Additionally, a definition of extracted characteristics is available. Using the HTTP, HTTPS, FTP, SSH, and email protocols, researchers constructed the abstract activities of 25 individuals for this dataset. Brute Force FTP, Brute Force SSH, DoS, Heartbleed, Web Attack, Infiltration, Botnet, and DDoS are some of the attacks used. They were executed on Tuesday, Wednesday, Thursday, and Friday in the morning and the afternoon. Sharafaldin et al. (2018). Database source: This data was generated by Sharafaldin et al. (2018). Specifically, the RF classifier, LR, and XGBoost have been examined among the ML systems in this study.
4.1 Experimental Goal
To the best of our knowledge, nobody has executed the following experiment,
  1. 1.
    Ensemble Stacking Classifier, where base classifiers are RF and XGBoost, the final LR classifier has never been implemented in SDN-based VANET as an IDS model. To gain optimum accuracy, FP/FN score and quick operation. And,
  2. 2.
    Bayesian Gaussian Hyperparameter Optimization to get the best parameter specs. And,
  3. 3.
    SelectKBest with chi² statistical test-based feature scoring contributed to this study’s best 10 feature selections.
  4. 4.
    The model’s performance was measured for the two datasets stated using Accuracy, Precision, Recall, F1, RMSE, and r2 Score, as well as the run-time of the IDS.
Finally, the above model will ensure the robustness and resilience of the proposed IDS model.
4.1.1 Virtual Machine Platform
To achieve the best possible hardware support for running the code, Microsoft Azure ML Studio Virtual Machine (VM) instance Standard E4s v3 (4 cores, 32 GB RAM, 64 GB HDD) and Processing unit CPU - Memory optimised were utilised for this experiment. The experiments for this research were carried out using Python 3.8.5, numpy Version 1.21.6, Pandas 1.1.5 and Scikit-Learn 0.23.1.
4.1.2 Supervised ML Approaches
Specifically, the RF classifier, LR, and XGBoost have been examined among the ML systems in this study.
Random Forest Classifier:
The randomisation must decrease correlation while retaining strength to increase accuracy. The forests under study grow each tree by applying inputs or combinations randomly chosen at each node. Accuracy results from the produced forests are favourably comparable to those of Adaboost. This class of methods has the following desirable qualities:
  • Its accuracy is comparable to Adaboost and occasionally much better.
  • It is relatively resistant to noise and outliers.
  • It moves more quickly than bagging or boosting.
  • It is straightforward, easily paralleled and provides helpful core estimates of error, strength, correlation, and variable relevance.
Logistic Regression:
LR is a linear model used to classify instead of predict. A logistic function is used for probabilities that describe the possible results of a single trial in this model. Typically, a sigmoid curve with the shape of a Sigmoid curve is called a logistic function or logistic curve. Wikipedia (2023)
XGBoost:
XGBoost is built using decision trees that have increased their gradients. Compared to other ML algorithms, XGBoost is fast and performs well. It is a technique for boosting machinery or doing machine boosting. Extreme Gradient Boosting, or XGBoost, helps maximise memory and hardware resources when used with tree-boosting techniques. The algorithm is enhanced, and the model is changed, which has benefits. The second derivative is used to perform Taylor expansion on the cost function, improving the results’ accuracy. The XGBoost method employs an additive training approach to optimise the objective function, suggesting that the outcome of the earlier stage affects the optimisation process in the subsequent stage. Gad et al. (2021)
4.2 Technical Implementation Methodology
This study will address multiclass problems using two approaches: Reducing Multiclass to Binary decision and multiclass methodology. The performances of the two approaches will then be discussed below.
Binary decision reduction:
Before feeding the data into ML algorithms, cleaning and preparation are crucial steps to achieve high performance. Unnecessary features may impact the performance of the chosen ML method. Missing values are regularly present in datasets.
To provide a relevant analysis, these values must be handled properly. After grouping the same attacks and splitting the datasets into two parts, training data and testing data, a Binary classifier was used to label the dependent columns in BENIGN and ATTACK. LabelEncoder is used to normalise labels so that the dataset only contains values between 0 and n classes-1. The model fitting and learning purposes are not similarly affected by variables measured at different scales and might lead to bias. Consequently, feature-wise normalisation, like the MinMax Scaling, is typically performed before model fitting to address this potential issue. Normalizer is applied row-by-row, while MinMaxScaler is done column-by-column. It balances all the information structures in the range [0, 1], otherwise in the range [-1, 1] in case there are negative values in the database.
General Multiclass methodology:
To begin this procedure, distinct datasets were constructed by keeping a specific attack and dropping the remainder of the attacks to create a new data combination for each particular attack vs BENIGN data.
Feature Selection:
The SelectKBest class from the scikit-learn library uses various statistical trials to select specific features. For example, the experiment used the chi-squared (chi2) statistical test as a scoring function for non-negative characteristics to pick the ten best features. The Chi-Square test is a statistical method for measuring the discrepancy between observed and projected data. This test can also be achieved to see if the data’s categorical variables are related. For example, it is helpful to determine whether a discrepancy between two categorical variables results from coincidence or an association between them. The Chi-Square Test is a method for quantifying the degree of disparity between expected and actual findings. To determine whether the results that were observed and those that were anticipated differed, the Chi-Squared statistic was used.
Hyperparameter Optimization: Bayesian Optimization with Gaussian Process:
BayesSearchCV incorporates "fit" and "scoring" methods. Using a cross-validated search of parameter settings, the parameters of the estimator used to apply these methods are optimised. In contrast to GridSearchCV, several parameter settings have been experimented with from the selected distributions rather than all parameter values. n_iter requires the number of parameter settings that are assessed. The ideal hyperparameters for the models can be found by efficiently scanning the hyperparameter space by utilising optimisation techniques throughout the design process of ML models. An estimator (a regressor or classifier) with its optimisation problem, a search space (configuration space), a search or enhancement method used to find hyperparameter combinations, and an evaluation function used to evaluate the performance of various hyperparameter setups. A hyper-domain parameter can be categorical, binary, discrete, continuous, or discrete. As a result, there are three types of hyper-parameters: continuous, discrete, and categorical. In real-world applications, continuous and discrete hyperparameters’ domains are typically constrained. On the other hand, conditionality can occasionally be seen in the hyper-parameter configuration space. Depending on the value, using or tuning a hyperparameter may be necessary. This study used hyperparameter optimisation to feed stacking base classifier RF.
Ensemble Stacking Classifier:
There are three different methods of ensemble learning: bagging, boosting, and stacking. The alternate voting methods are bagging and boosting. Standardised models are used in the bagging strategy to estimate the test data class. Predictions from homogeneity chosen models are initially recorded. The test data are then assigned to the class that most models predict. Similar to boosting, the training phase of the models includes extensive training for the misclassified data. The model with the highest accuracy is finally chosen as the classifier for the test data. Because stacking is a generic framework with numerous ensemble approaches, it is a successful strategy. Base and meta-learning are two of the learning layers. A training data set is used to teach the initial (base) learners in base learning. After completion of training, the base learners generate a new dataset for the meta-learner. The meta-learner is then trained using a clean set of training data. Finally, the test set is categorised using the trained meta-learner. The ideal foundation learner should be chosen as it is an essential stacking component. In other words, numerous base learners should be used for the training data set rather than just one. The primary distinction between stacked and other ensemble approach methods is the application of meta-level learning-based classification as the final classification in stacking. As mentioned, this study utilised random forest (RF) and XGboost as base learners in a stacked ensemble model. Lastly, the final learner is logistic regression.
Experimantal setup:
Using the provided VeReMi and CICIDS2017 datasets, the intrusion detection system adopts ensemble learning classifiers to tackle issues linked to multiple classification models. The CICIDS2017 set presents an enormous dataset based on many attacks, including DDoS and PortScans, recorded for numerous days. The VeReMi set contains a few comma-separated Values(CSV) files encapsulating acts of assault by a specific aggressor. The organisation phase required eliminating any records with null values denoted as NaNs or Infinities —to ensure the actual credibility of the models formed. The VeReMi labels are then transformed from numerical codifications into substantial descriptions, such as "Constant Attack" or "Random Attack.". The CICIDS2017 CSV file deduces that some attack types are formed into unique invasion types with a broader category. The viably divided datasets training and validation data sets at an 80:20 ratio to grant every variety proportion in training modelling. Normalisation by scaling methods like MinMaxScaler ensures the balance of size magnitudes of integer inputs into the numerical system to ensure the model is incapacitated to give visual bias. A blend of models like RandomForest and XGBoost has side benefits with Logistic Regression as the second phase model. For attacks less regularly known, like Heartbleed in VeReMil, the count of cross-validation regression folds is lowered to CV = 4 to prevent overfitting against the undersized represents of points. Metrics including accuracy, precision, recall, F1 score, and more sophisticated measurements like RMSE (Root-mean-square deviation) were used to assess the model.
5 Results and Evaluation
In the suggested model, accuracy is the most significant quality indicator. This precision is used to evaluate the model’s performance in the intrusion detection system. The performance of the suggested system has been assessed using accuracy (ACC), detection rate (DR), precision, and recall metrics. True positive (TP), true negative (TN), false positive (FP), and false negative (FN), accordingly.
During anomaly identification, the training and testing phases utilise distinct datasets. However, as evidenced by the results, the accuracy percentage is exceptionally high and can be considered outstanding. In addition, the other performance metric exhibits encouraging results.
Table 4: Estimators for Multiclass Approach.
Datasets Multiclass Classification
CICIDS2017
estimators = [(’rf’, RandomForestClassifier(n_estimators= 26, max_depth= 50,
min_samples_split= 2,
min_samples_leaf= 1, criterion= ’entropy’)),
(’xgb’, xgb.XGBClassifier(eval_metric=’rmse’, use_label_encoder=False))]
clf = StackingClassifier(estimators=estimators,
final_estimator=LogisticRegression(solver=’lbfgs’, max_iter=800000))
VeReMi
estimators = [(’rf’, RandomForestClassifier(n_estimators= 100, max_depth= 50,
min_samples_split= 10,
min_samples_leaf= 2, criterion= ’entropy’)),
(’xgb’, xgb.XGBClassifier(eval_metric=’rmse’, use_label_encoder=False))]
clf = StackingClassifier(estimators=estimators,
final_estimator=LogisticRegression(solver=’lbfgs’, max_iter=800000),
stack_method=’auto’, passthrough=False, n_jobs=-1)
Table 5: Estimators for Multiclass Approach for small attack size.
Attacks CICIDS2017
Heartbleed
Infiltration
estimators = [(’rf’, RandomForestClassifier(n_estimators= 1, max_depth= 5, random_state=None,
min_samples_split= 11, min_samples_leaf= 1, criterion= ’entropy’)),
(’xgb’, xgb.XGBClassifier(eval_metric=’rmse’, use_label_encoder=False))]
clf = StackingClassifier(cv=4, estimators=estimators, final_estimator=LogisticRegression
(solver=’lbfgs’, max_iter=800000), stack_method=’auto’,
passthrough=False, n_jobs=-1)
Table 6: Model performance evaluation table for Binary Classification Approach
Datasets Accuracy Precision Recall F1 RMSE r2 Score Time (hh:mm:ss)
CICIDS2017 99.92% 99.75% 99.87% 99.81% 0.027 0.99 00:27:37
VeReMi 99.68% 99.60% 99.31% 99.45% 0.05 0.98 00:07:48
Table 7: Model performance evaluation using Multiclass Approach for CICIDS2017 dataset.
Attacks Accuracy Precision Recall F1 FN FP TN TP RMSE R2 Time
DoS 99.984 99.857 99.988 99.922 6 72 454193 50336 0.012 0.998 01:59:46
PortScan 99.958 99.386 99.987 99.686 4 196 454068 31757 0.020 0.993 01:24:27
DDoS 99.997 99.972 99.988 99.980 3 7 454257 25602 0.004 0.999 01:44:54
Brute Force 99.998 98.866 100 99.429 0 5 454259 436 0.003 0.988 01:31:02
Web Attack 99.998 98.863 99.770 99.315 1 5 454259 435 0.003 0.986 01:20:33
Bot 99.961 70.185 96.930 81.417 12 161 454104 379 0.019 0.557 01:37:20
Infiltration 100 100 100 100 0 0 454265 7 0 1 00:57:12
Heartbleed 99.999 80 100 88.888 0 1 908528 4 0.001 0.749 00:16:59
Table 8: Model performance evaluation using Multiclass Approach for VeReMi dataset.
Attacks Accuracy Precision Recall F1 FN FP TN TP RMSE R2 Time
Eventual Stop 99.705 99.355 98.065 98.706 219 72 87414 11100 0.054 0.970 00:09:42
Random 99.994 100 99.918 99.959 5 0 87486 6097 0.007 0.999 00:09:23
Constant Offset 99.989 100 99.835 99.917 10 0 87486 6085 0.010 0.998 00:08:21
Constant 100 100 100 100 0 0 87486 6095 0 1 00:02:10
Random Offset 99.962 99.897 99.507 99.702 29 6 87480 5863 0.019 0.993 00:13:32
Root-mean-square error (RMSE) is a regularly utilised measure of the changes between the expected values (sample or population values) and the actual values. The RMSE is the square root of the second instance of the deviations between expected and actual values or the quadratic mean of these deviations. When these deviations are computed across the test dataset used for the assessment, they are known as residuals; when estimated out-of-sample, they are known as errors (or prediction errors). The RMSE integrates the magnitudes of errors in forecasts for different data points into a familiar figure of projecting skill. Due to its scale dependence, RMSE is a measure of accuracy used to assess the prediction mistakes of various models for a single dataset and not between datasets. RMSE is always positive, and a value of 0 indicates that the model fits the data perfectly (which is practically never attained in practice). A smaller RMSE is often preferred over a greater one. However, because the measure depends on the scale of the employed numbers, comparisons between distinct data kinds are erroneous. The effect of each error on RMSE is equal to the error’s squared size; thus, more significant errors have a disproportionate effect on RMSE.
  • RMSE is the common deviation of the residuals.
  • RMSE shows average model prediction fault.
  • The lower values specify a good fit.
  • It is measured in the same units as the Target variable.
The percentage of the variation in the dependent variable that can be predicted from the independent variable is known as the coefficient of determination in statistics. It is symbolised by the symbols $R2$ or $r2$. It is a statistic applied to statistical models whose main objective is to either predict future outcomes or test hypotheses using data from other relevant sources. Based on the percentage of the overall variation in outcomes that the model is responsible for explaining, it gauges how effectively the model mirrors observed results. In most cases, the coefficient of determination lies between 0 and 1.
5.1 Experimental results
CICIDS2017 and VeReMi datasets are used to evaluate the performance of binary classification models and multiclass classification models. The binary classification method achieved good accuracy, precision, recall, and F1 score for both datasets. The model scored 99.92% accuracy, 99.75% precision, 99.87% recall, and 99.81% F1-score for CICIDS2017. The model scored 99.68% accuracy, 99.60% precision, 99.31% recall, and 99.45% F1-score for VeReMi. The RMSE and r2 values show that the model’s predictions were correct and fit the data well. The model trained on CICIDS2017 took longer than the model trained on VeReMi.
The multiclass classification strategy has been implemented using the CICIDS2017 and VeReMi datasets. The models were taught to recognise eight different types of attacks. For CICIDS2017, the model achieved high accuracy, precision, recall, and F1 score on all attacks. The model’s overall accuracy was 99.99%, and the lowest accuracy for an attack was 99.96% for PortScan. The model for VeReMi achieved great accuracy, precision, recall, and F1 score for all five types of attacks. The model’s overall accuracy of the model was 99.99%, and the lowest accuracy achieved for an attack was 99.71% for Eventual Stop.
The RMSE and r2 values indicate that the model’s predictions were correct and fit the data well. On the other hand, the model trained on the Bot attack for CICIDS2017 had a low r2 score of 0.557, indicating that the model’s predictions were not well-fitted to the data and the models’ training time varied according to the number of attacks to be classified. The CICIDS2017 model took longer, from 00:57:12 to 01:59:46, while the VeReMi model took between 00:02:10 and 00:13:32.
FN and FP are critical indicators of a classification model’s success since they accurately identify positive and negative situations. A successful model should have low FN and FP rates, indicating that it can detect positive and negative cases correctly. This model achieved this with an excellent score for multiclass classifications. To be more specific, this model performed effectively even though infiltration and heartbleed attack entries in CICIDS2017 were relatively shallow. The estimator table demonstrates how parameters were applied to produce the best results.
(a) Classification Report for CICIDS2017
(b) Classification Report for VeReMi
Figure 4: Classification Report by attacks.
These slight variances in performance for attack detection in the IDS model, using the CICIDS2017 and VeReMi datasets, can be enlightened better based on the model’s precision, recall, F1 scores, false positives, and false negatives. In the CICIDS2017 dataset, precision and F1 scores are significantly lower than for other attacks such as DoS and DDoS. This lower performance indicator might suggest that Bot attacks display patterns so akin to behaviour that the model finds it challenging to detect them without a false hit on benign activities as an attack. Therefore, the high number of positives for Bot attacks indicates that the model is more likely to misclassify traffic as malicious due to similarities to Bot-like behaviours. However, even though Heartbleed has the highest true positives and accuracy, the precision performance indicator is second-to-last in the dataset. The reduced precision of this particular attack indicates a Heartbleed in the dataset, and the model’s confidence is undermined by its experience with such examples. Alternatively, features of Heartbleed attacks may overlap significantly with behaviours, posing classification challenges. As highlighted in the dataset, except for the Eventual Stop attack, the detection performance is crucial for attacks, presenting a slightly lower precision and higher false negatives, implying that it is challenging for the model to detect this attack pattern consistently, perhaps because the specific characteristics of the Eventual Stop attack are not as straightforward or consistent. Besides, the high precision score for Random and Constant Offset attacks indicates that these attacks have consistent patterns, which the model can learn quickly. These might have features that differ significantly from behaviour. Hence, they are easily identified by the model. Moreover, the false negative rates will indicate a balance between specificity and sensitivity in the model. High false negative rates, as observed in the Eventual Stop attack, suggest that the model is precise in its predictions but is not sensitive enough to detect this attack in some cases. Random Forest and XGBoost as base learners and Logistic Regression as a learner have shown model effectiveness, considering now the performance evaluation for CICIDS2017 and VeReMi datasets. The model implementation involved a combination of these algorithms purposely to use their strengths and mitigate their weaknesses through stacking. The variance in model effectiveness against types of attacks is influenced by the complexity and frequency of attack patterns, the degree of similarity among harmful and non-harmful features, and the data volume and quality that have been used to train the model. These factors contribute to how efficiently the model has learned, ultimately influencing its accuracy in detecting and classifying types of attacks.
In the end, in the CICIDS2017 and VeReMi datasets, binary and multiclass classification models performed well in identifying attacks. The models’ high accuracy, precision, recall, and F1-score and the low RMSE and high r2 values show that the model’s predictions were accurate and well-fitted for the data. The training duration varies according to the size of the dataset and the number of attacks to be classified.
Moreover, to further solidify the claim that the proposed approach was robust and efficient, we compared our stacked ensemble model with other methods, as presented in Table 3. For instance, Gad et al. (2021) attained 99.1% accuracy while employing XGBoost after feature selection, Türkoğlu et al. (2022) obtained 99.35% accuracy with Decision Tree after MRMR feature selection, Zeng et al. (2019) used DeepVCM with CNN and LSTM to realise 98.5% accuracy, and Aneja et al. (2018) developed an ANN with a genetic algorithm to get 99% accuracy. Comparatively, our proposed stacked ensemble model introduced higher accuracy levels at 99.92% with the CICIDS2017 dataset and 99.68% with the VeReMi dataset.
5.2 Challenges and Experiences to Implementation
Depending on database size and hardware capacity, the elapsed time is a big challenge in the classification phase. Hardware capacity can be upgraded to get quick results, but it will not be cost-effective for customers. The two publicly available datasets, VeReMi and CICIDS2017, are the primary source of our study’s glaring errors. Although these synthetic datasets are widely used in research, they cannot fully represent the intricacy of actual vehicle networks. More precisely, other notable attack types, such as denial of service or time-sensitive threats, are not included in the VeReMi dataset. It primarily focuses on misbehaviour attacks, such as fabricating location data. Due to the constraint mentioned above, our model may be unable to identify as many harmful activities as possible in real-world scenarios.
Developing a stacked ensemble is a proposed learning method that uses Random Forest, XGBoost, and Logistic Regression algorithms to obtain a reliable Intrusion Detection System for Software-Defined VANETs. Random Forest and XGBoost are computationally expensive during training because they have many trees and extensive data sorting. In comparison, Logistic Regression is simple but requires more iterations to converge. The storage requirement pertains to the number of trees and storage units and their weight matrices in Logistic Regression. The complexity of the ensemble model will affect the individual classifiers’ training and storage space. The model is efficient and effective as an IDS. The approach is practical in all aspects. It has enhanced the various trade-offs in achieving high accuracy and higher performance in terms of efficiency, rendering it appropriate for the operation of the dynamic and resource-constrained network, which are the VANETs. Generally, from the findings, VANETs have various requirements and windows for realising the effectiveness of improving an IDS.
5.3 General Implementation method in SDN-based VANET
The SDN controller has two unique interfaces, one on the northbound side that communicates with the cloud, orchestration, and management layer and the other on the southbound side that talks to the data plane layer, as shown in Figure 1. In most of the SDN controllers that are now available, the northbound interface is standardised as REST APIs. Only the vendor’s implementation of the southbound interface is supported. Some use OpenFlow, BGP, Netconf, and other open standards-based protocols, while others use proprietary interfaces. An SDN controller is a software scheme or group of systems that collectively offers the following:
(A) A database will be used to manage the state of the network and, in some situations, to distribute this state. These databases act as a store for data obtained by the managed network components and associated software, as well as data managed by SDN applications (including network state, learned topology, specific ephemeral arrangement data, and control session data). In some circumstances, the controller could have several targeted data procedures. Different in-memory database techniques may also be used in other situations.
(B) A large data type depicts the connections between controlled resources, rules, and other services the controller offers.
(C) An application can access the controller services through a modern, frequently RESTful (Representational State Transfer) application programming interface (API). This makes the vast majority of controller-to-application interactions more accessible. The database schema representing the controller’s services and features should be used to construct this interface. In certain situations, an application framework that generates API code from a model contains both the controller and its API. Some systems offer sophisticated DevOps (software development and operations) that permit the development of core capabilities and the publishing of APIs for new modules, including those that facilitate the continuous expansion of controller abilities.
(D)Between the controller and the connected agents in the network elements, a secure TCP control session
(E) A protocol that adheres to standards for the delivery of application-driven network state on network components
(F) A path computation system, a device, topology, service discovery method, and possibly additional network- or resource-centric information services.
The baseline is this IDS solution can be installed in the OpenFlow switch in the control plane layer.
Advantages and Limitations of the proposed method
Advantages:
  • Both models are a good option when a single classifier cannot predict with optimum accuracy.
  • Three-stage ensemble method will reduce the risk of Inference Attacks. Salem et al. (2018)
  • Quick run-time with impressive accuracy.
  • Even performs well against small-scale attacks.
Limitation:
It cannot detect encrypted and unknown attacks.
6 Conclusion
In conclusion, this research merely targeted an effective solution to detecting a cyber-attack in SDN-based VANETs, a vital network vulnerability. Therefore, even after taking note of the limitation of the dataset, the proposed ensemble-based ML model using Random Forest (RF) in combination with XGBoost and Logistic Regression has adequate accuracy in attack detection and the ability to predict with low false positives. This demonstrates that the model has a higher accuracy rate and level than the other available models, specifically in relation to the NN. Besides the above considerations, many deep learning approaches are costly and time-consuming. The proposed method presents a faster solution to detecting an attack on autonomous vehicles. Therefore, we recommend this solution be applied within an SDN controller for VANETs to enhance network performance and monitoring in the intelligent network design. Banitalebi Dehkordi et al. (2021) Furthermore, more significant hyperparameter tuning could achieve even higher accuracy for potential total accuracy.
Future Scope
Research is an endless endeavour. The future potential of this thesis resides in its ability to identify attacks using encrypted malicious data. Finally, proceed with global detection and deep learning prospects. This concept would pioneer the development of a deployable IDS solution for SDN-based VANETs. Moreover, it is possible to implement this model in other public datasets or make an extensive public synthetic VANET database with known attacks to make the model more universal and test it in the simulator before conducting a field demonstration.
The main advantage of the ensemble learning IDS is the fact that several algorithms accustomed to detecting noticeable network abnormalities are combined. One may consider the need for an algorithm to detect time-related abnormalities in the complications of DoS assailment to VANETs; others may indicate appropriate spatial irregularities that imply location spoofing. Finally, a model requires a dynamic updating mechanism, such as self-learning intrusion detection, that can adjust to new categories of threats. This constraint may be met by incorporating real-time databases and continuous learning guidelines. One may also revise future recognised assault scenarios with more sophisticated machine-learning strategies. Bonafide improves our discovery power by adding various attributes to the set. The system’s versatility is crucial to its operation, and a flexible framework that can render decoding algorithms promptly and encompasses a hazard definition is needed. The study would be substantial for establishing cybersecurity policies as it provides a model to be used for many purposes.
Declarations
Competing interests
The authors declare that they have no competing interests.
Authors’ contributions
Shakil Ibne Ahsan led the research with guidance and research support from other authors as part of research conducted at the University of the West of England, UK. All authors contributed to the preparation of the manuscript.
Availability of data and materials
Not applicable.
Funding
Not applicable.
Acknowledgements
Thank you to the School of Computing and Creative Technologies at the University of the West of England for providing the necessary support to enrich this study. The authors would also like to thank the reviewers for their valuable feedback, which has helped strengthen the article through peer review.